RC meeting and Run Plan, Monday February 27th

From SBS wiki
Jump to: navigation, search

RC Report


  • Completed 45 C on Fringe
  • Completed Successful Moller

Recent Issues

  • None

Current Status

  • Target Studies

Run Plan

  • After this complete Optics
  • Prep Hall for Target Replacement

Data Collection and Polarization Status

  • 45.027 C of uncorrected charge for this kinematic (52.707 C of corrected charge which is 33% of the plan)
  • Last NMR 46.9%

Other Topics/Status Updates

Requested Accesses to the Hall

Don Jones Jessie or Techs?